OpenStudy (anonymous):

A line passes through (2, 8) and (4, 12). Which equation best represents the line? y = 1 over 2 x + 6 y = 2x + 7 y = 2x + 4 y = 1 over 2 x + 10

OpenStudy (astrophysics):

Well first find the slope, using the following formula: \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] where your coordinates represent \[(x_1,y_1)~~\text{and}~~(x_2,y_2)\] Once you manage to get that, plug in m and your first coordinates in the point - slope form: \[y-y _{1}=m(x-x _{1})\] do some little algebra and eventually you will have slope - intercept form \[y=mx+b\]

OpenStudy (ehsan18):

let m be the slope m=(y2-y1)/(x2-x1) m=(12-8)/(4-2) m=2 Now Eq. of line is y-y1=m(x-x1) y-8=2(x-2) y=2x-4+8 y=2x+4 is the answer.